just want the steps of these problem (x+3)/(x-2)=(x+4)/(x-3),thanks
Answer by AnlytcPhil(1116)

 (x+3)/(x-2)=(x+4)/(x-3),
 
 x+3     x+4 
————— = —————
 x-2     x-3

This is a proportion, that is, of the form

FRACTION = FRACTION

When there are no common factor of the denominator, 
cross-multiply:

RULE for CROSS-MULTIPLYING:

(UPPER LEFT)·(LOWER RIGHT) = (LOWER LEFT)·(UPPER RIGHT)

Using the CROSS-MULTIPLICATION rule:

 x+3     x+4 
————— = —————
 x-2     x-3

(x+3)·(x-3) = (x-2)(x+4)

FOIL out both sides:

x2 - 3x + 3x - 9 = x2 + 4x - 2x - 8

          x2 - 9 = x2 + 2x - 8

Subtract x2 from both sides

             -9 = 2x - 8

Add 8 to both sides

             -1 = 2x

Divide both side by 2

             -1     2x
            ———— = ————
              2      2

Cancel the 2's on the right:
                    1 
             -1     2x
            ———— = ————
              2      2
                     1

            -1/2 = x

Edwin
AnlytcPhil@aol.com